### v2

@@ 3.2.1 Feature binarization
In spite of information loss, input binarization is a direct way to denoise and avoid overfitting. We perform binarization on path-signature-extracted features and then use them as input data of the dual-channel autoencoder network for further learnable feature compression. Specifically, given subjects $i$ and $j$ and their concatenated features, $n_{i}$ and $n_{j}$, we have the input pair $F_{i,j}^{cat}=\{n_{i},n_{j}\}$. For each feature dimension, we extract the median as the binarization threshold. Feature values larger than the median are set to 1, and others are set to 0. After binarization, the data pair is denoted as $F_{i,j}^{bin}=\{b_{i},b_{j}\}$. During this step, the noises are smoothed out. We will show in the experiments that we achieve better performance with binarization.
@@ 3.2.2 Dual channel unsupervised feature compressor
9:              Forward $P_{i,j}^{q-1}$ on $AE_{1,2}^{q}$ batch-wise, obtaining the result $P_{i,j}^{q}$ and retain the gradients.
@@
11:         Here obtaining $P_{i,j}^{k}$, which is the compressed feature vector.
@@
14:              Forward $P_{i,j}^{q-1}$ on $AE_{1,2}^{q}$ batch-wise, obtaining the result $P_{i,j}^{q}$ and retain the gradients.
@@ 4.2 Ablation experiments
In this section, we conduct ablation experiments to evaluate the effectiveness of our model. In order to prevent over-fitting of unsupervised training, the concatenated feature is binarized before entering the dual-channel autoencoder network. The ablation experiment results shown in Table 1 can prove that the performance of the algorithm after binarization is better, where the autistic recalling ability of the model is decreased without binarization. This result may indicate that the key structural and developmental information for autism diagnosis is relatively invariant during binarization.
Table 1: The model effect with & without binarization operation.
@@
Table 2: The model effect with different brain region weight constraint mechanisms.
@@
Ablation experiments of each module are also carried out to verify the corresponding module effectiveness, with results shown in Table. 3. “Our’s” denotes our algorithm, which is deploying all modules. “W/o weight & AE” denotes the algorithm obtained on the basis of the Siamese verification network with PS feature as Zhang et al. (2020), where neither dual-channel autoencoder modules nor brain region weight constraint mechanism modules is used. In a similar way, “W/o weight” denotes the algorithm after removing the brain region weight constraint module in the test phase. “W/o PS” denotes an algorithm that does not add path signature features. “W/o AE” denotes an algorithm that does not add the dual-channel autoencoder module.
@@
Table 3: Ablation experiments of each module of our model.
@@ 4.3 Comparison with other methods
Finally, we compare our own algorithm with other related works, as shown in Table 4. SVM denotes the Support Vector Machine (SVM) classification method. RF denotes the random forest method. SIAM_PS is the work of Zhang et al. (2020), and DL+SVM refers to the work of Hazlett et al. (2017). “Our’s” denotes our algorithm. Since the imbalance in the number of data categories is a major characteristic of the dataset, the compared algorithms generally have poor recall rates, which means the corresponding sensitivity value is very low in Table 4. The DL+SVM method have even 0 recall with the categorically imbalanced dataset. By comparing our results with other algorithms, we can find that our model is specifically effective for the recall of autism samples, which increases the model sensitivity while maintaining high accuracy, represented by a high F1-score.
Table 4: Comparison with other autism diagnosis models under few-shot and class-imbalanced conditions.
@@
Our’s 0.585 0.829 0.633 0.874
@@ 4.4 Cortical region importance analysis
Our input features of the model are based on the infant dedicated cortical surface atlas proposed by Wu et al. (2019), from which we used features of 70 cortical ROIs. The importance of each brain region is calculated with the method described in section 3.5. Fig. 6 presents the visualization of the most influential cortical regions during the classification.
@@ 5 Discussion and conclusions
Data used in this paper can be shared via a formal data sharing agreement with us.
@@ 6 Author contributions statement
Zhuowen Yin (Conceptualization, Formal Analysis, Investigation, Methodology, Resources, Software, Validation, Visualization, Writing – original draft, Writing – review & editing), Xinyao Ding (Conceptualization, Methodology, Software, Writing – original draft), Xin Zhang (Conceptualization, Methodology, Project administration, Resources, Supervision, Writing – review & editing), Zhengwang Wu (Data curation, Writing – review & editing), Li Wang (Data curation, Resources), and Gang Li (Conceptualization, Data curation, Project administration, Resources, Supervision, Writing – review & editing).

### v3

@@ 3.2.1 Feature binarization
In spite of information loss, input binarization is a direct way to denoise and avoid overfitting. The idea of input feature binarization was first introduced by Rastegari et al. (2016), where the authors have showed that input binarization is an effective preprocessing step that only brings minor performance decrease. Feature binarization is then further used in small sample size learning scenarios, like in Hazlett et al. (2017). The rationale of doing so is to avoid overfitting on small datasets by reducing data complexity, as discussed in Bejani and Ghatee (2021). Therefore, we perform binarization on path-signature-extracted features and then use them as input data of the dual-channel autoencoder network for further learnable feature compression. Specifically, given subjects $i$ and $j$ and their concatenated features, $n_{i}$ and $n_{j}$, we have the input pair $F_{i,j}^{cat}=\{n_{i},n_{j}\}$. For each feature dimension, we extract the median as the binarization threshold. Feature values larger than the median are set to 1, and others are set to 0. After binarization, the data pair is denoted as $F_{i,j}^{bin}=\{b_{i},b_{j}\}$. During this step, the noises are smoothed out. We will show in the experiments that we achieve better performance with binarization.
@@ 3.2.2 Dual channel unsupervised feature compressor
9:              Forward $F_{i,j}^{q-1}$ on $AE_{1,2}^{q}$ batch-wise, obtaining the result $F_{i,j}^{q}$ and retain the gradients.
@@
11:         Here obtaining $F_{i,j}^{k}$, which is the compressed feature vector.
@@
14:              Forward $F_{i,j}^{q-1}$ on $AE_{1,2}^{q}$ batch-wise, obtaining the result $F_{i,j}^{q}$ and retain the gradients.
@@ 4.2 Ablation experiments
In this section, we conduct ablation experiments to evaluate the effectiveness of our model. In order to prevent over-fitting of unsupervised training, the concatenated feature is binarized before entering the dual-channel autoencoder network. The ablation experiment results shown in Table 1 compare the model’s performance when using the binarized features versus the original normalized features. It can prove that the performance of the algorithm after binarization is better, where the autistic recalling ability of the model is decreased without binarization. This result may indicate that the key structural and developmental information for autism diagnosis is relatively invariant during binarization.
Table 1: The model effect with & without binarization operation, evaluated with 10-fold cross-validation.
@@
Table 2: The model effect with different brain region weight constraint mechanisms, evaluated with 10-fold cross-validation.
@@
Ablation experiments of each module are also carried out to verify the corresponding module effectiveness, with results shown in Table. 3. “Ours” denotes our algorithm, which is deploying all modules. “W/o weight & AE” denotes the algorithm obtained on the basis of the Siamese verification network with PS feature as Zhang et al. (2020), where neither dual-channel autoencoder modules nor brain region weight constraint mechanism modules is used. In a similar way, “W/o weight” denotes the algorithm after removing the brain region weight constraint module in the test phase. “W/o PS” denotes an algorithm that does not add path signature features. “W/o AE” denotes an algorithm that does not add the dual-channel autoencoder module.
@@
Table 3: Ablation experiments of each module of our model with 10-fold cross-validation.
@@ 4.3 Comparison with other methods
Finally, we compare our own algorithm with other related works, as shown in Table 4. SVM denotes the Support Vector Machine (SVM) classification method. RF denotes the random forest method. SIAM_PS is the work of Zhang et al. (2020), and DL+SVM refers to the work of Hazlett et al. (2017). “Our’s” denotes our algorithm. All the comparison experiments are conducted under the same conditions on the same dataset, as introduced in the Materials section. Since the imbalance in the number of data categories is a major characteristic of the dataset, the compared algorithms generally have poor recall rates, which means the corresponding sensitivity value is very low in Table 4. The DL+SVM method have even 0 recall with the categorically imbalanced dataset. By comparing our results with other algorithms, we can find that our model is specifically effective for the recall of autism samples, which increases the model sensitivity while maintaining high accuracy, represented by a high F1-score. We also include Fig. 6 to better present the statistical property of our comparison experiment results visually. It is shown that, although our results have some fluctuations due to small sample size, standard deviations are in reasonable ranges.
Table 4: Comparison with other autism diagnosis models under few-shot and class-imbalanced conditions, on the same dataset with 10-fold cross-validation.
@@
Ours 0.585 0.829 0.633 0.874
[…]
Figure 6: Visualizing the statistical results of accuracies of different methods. Error bars indicate standard deviations.
Additionally, we summarize some recent works using MRI data for autism classification in Table 5, to provide a broader view on the progress of the field. Of note, those various works used very different input data types and datasets with different age ranges, so this is not a quantitively fair comparison, but only an additional general review. Works listed in the table utilized different imaging modalities and computational methods for autism diagnosis. Compared with other works, our method performs well under practical scenarios of scarce, imbalanced, infantile, and sMRI-only data.
Table 5: A brief summary of recent works on autism classification with MRI data. The short forms used in the table are: ASD, Autism Spectrum Disorder; DL, Deep Learning; DTI, Diffusion Tensor Imaging; FCM, Functional Connectivity Matrices; MBN, Morphological Brain Networks; NC, Normal Controls; RF, Random Forests; SVM, Support Vector Machines.
[…]
Zhao et al. (2018) 100 (54/46) fMRI multi-order FCMs under 15 0.83 0.81 0.82 0.80
Soussia and Rekik (2018) 341 (155/186) sMRI multi-order MBNs $\sim$ 17 - 0.6170 - -
Zhang et al. (2018) 149 (70/79) DTI SVM 6-18 - 0.7833 0.8481 0.7286
Sen et al. (2018) $1$ (538/573) fMRI+sMRI DL+SVM 7-64 - 0.6431 0.6832 0.6000
Li et al. (2018) 264 (55/209) sMRI DL 2 - 0.7924 - -
Kong et al. (2019) 183 (78/104) sMRI DL $\sim$ 15 - 0.9039 0.8437 0.9588
Rakić et al. (2020) 817 (368/449) fMRI+sMRI DL 7-64 - 0.8445 0.80 0.88
Kim et al. (2022) 106 (58/48) sMRI+DTI RF 3-6 - 0.888 0.930 0.838
Gao et al. (2022) 1994 (N/A) sMRI DL 5-64 - 0.6785 0.6166 0.7336
Ours 157 (30/127) sMRI DL 0-1 0.585 0.829 0.633 0.874
@@ 4.4 Cortical region importance analysis
Our input features of the model are based on the infant dedicated cortical surface atlas proposed by Wu et al. (2019), from which we used features of 70 cortical ROIs. The importance of each brain region is calculated with the method described in the Region importance extraction part of the Method section, and the importance is normalized into the range of 0.2 to 1 for better visualization. Fig. 7 presents the visualization of the most influential cortical regions during the classification.
@@
We also analyzed and included the importance of the gender features following the same extraction and normalization method in the Region importance extraction part of the Method section, and compared them with brain regional features. The normalized gender feature importance is 0.170, which is less than all the brain region importance weights ranging from 0.2 to 1, indicating that our model is paying less attention to gender information, compared to each of the brain regional MRI features.
However, since our gender feature input is 2-dimensional and the MRI features of each brain region are expanded to 18-dimensional, following the feature extraction process described in the Methods section, the importance comparison between gender and brain regional features may be biased even after normalization. So we additionally conducted ablation experiments with and without gender information as an input feature. For the “w/o gender info” experiments, we remove the gender features in the input and keep all other features. The ablation experiment results are shown in Table 6 below. The results showed that removing gender information negatively affects the model performance. This ablation experiment indicates that gender information is a helpful supplement to image data in autism diagnosis.
Table 6: The model performance with and without gender information, evaluated with 10-fold validation.
[…]
w/o gender info 0.484 0.797 0.500 0.866
With gender info (Ours) 0.585 0.829 0.633 0.874
[…]
@@ 5 Discussion and conclusions
Data used in the preparation of this manuscript were obtained from the NIH-supported National Database for Autism Research (NDAR). NDAR is a collaborative informatics system created by the National Institutes of Health to provide a national resource to support and accelerate research in autism. This manuscript reflects the views of the authors and may not reflect the opinions or views of the NIH or of the submitters submitting original data to NDAR.
@@ 6 Author contributions statement
Zhuowen Yin (Conceptualization, Formal Analysis, Investigation, Methodology, Resources, Software, Validation, Visualization, Writing – original draft, Writing – review & editing), Xinyao Ding (Conceptualization, Methodology, Software, Writing – original draft), Xin Zhang (Conceptualization, Methodology, Project administration, Funding acquisition, Resources, Supervision, Writing – review & editing), Zhengwang Wu (Data curation, Writing – review & editing), Li Wang (Data curation, Resources), Xiangmin Xu (Conceptualization, Project administration, Funding acquisition, Resources, Supervision), and Gang Li (Conceptualization, Data curation, Project administration, Resources, Supervision, Writing – review & editing).
7 Acknowledgments
This work was supported by the Guangdong Provincial Key Laboratory of Human Digital Twin (2022B1212010004 to X. Z. and X. X.), and the Fundamental Research Funds for the Central Universities (No. 2022ZYGXZR104 to X. Z.).